# SECURE: Benchmarking Generative Large Language Models for Cybersecurity Advisory

### v1

@@ 3.2.1. Extraction Task
3.2.2. Understanding Task
Our second task dataset is designed to evaluate the ability of LLM to comprehend and understand the security-related text. Given the continuously evolving landscape of cybersecurity, it is crucial that LLMs can assimilate new information and use it to generate accurate responses. For this purpose, we utilize the CVEs (Common Vulnerabilities and Exposures) published in 2024, available at CVE project repository (Project, 2024a). CVE is better for this task since they are always updated but CWEs are more appropriate for the extraction task as they are updated less frequently.
To ensure the integrity of our evaluation, we verified that none of the pretrained models we selected for evaluation (ChatGPT-4 (gpt, 2024b), ChatGPT-3.5 (gpt, 2024a), Llama3-70b (lla, 2024a), Llama3-8b (lla, 2024b), Gemini-Pro (gem, 2024), Mistral-7B (Jiang et al., 2023) , Mixtral-8x7b (mix, 2024)) had access to these CVEs during their training phase by confirming their training-cut-off date (See Table 5). This precaution guarantees that the models have not been exposed to this specific information and thus must rely on their comprehension abilities, and therefore, maintain the validity and reliability of SECURE.
@@ 3.2.2. Understanding Task
Table 2. An example row from RERT dataset. Overview & vulnerability description shortened for readability.
@@ 3.2.1 Path Traversal CWE-22 There are multiple ways in LAquis
3.2.3. Reasoning Task
Our third task evaluates the LLM’s reasoning capability in the context of ICS security through the Risk Evaluation Reasoning Task (RERT). To create RERT, we compile a risk assessment dataset using cybersecurity advisories from the Cybersecurity and Infrastructure Security Agency (CISA) (Advisories, 2024). We focus on all significant ICS-related advisory reports, each containing an Executive Summary, Risk Evaluation, Technical Details (including Vulnerability Overview, Affected Products, Background Researcher), and Mitigations. Reports lacking essential information are discarded to ensure data quality, and the remaining comprehensive reports are processed for inclusion.
@@ 3.2.3. Reasoning Task
In addition, we create another dataset CPST (CVSS Problem Solving Task) to measure the problem-solving skills of LLMs in security. Specifically, we use the CVSS (Common Vulnerability Scoring System) framework and manually collect 100 unique CVSS3.1 vector strings from the Cybersecurity and Infrastructure Security Agency (CISA) (Advisories, 2024). These scores, in the range of 0-10, can be computed using the CVSS calculator which uses the Base, Temporal, and Environmental scores to determine the overall severity of a vulnerability (cvs, 2024). This task involves using an existing formula from CVSS3.1 standard and computing a value. This computation allows the evaluation of the problem-solving skills of LLMs in practical security settings.
3.3. Dataset Validation
@@
Our analysis revealed several categories of issues within the generated questions. Some questions were deemed unanswerable from the provided context, particularly within the True/False format. In the Multiple Choice Question (MCQ) format, we encountered instances where multiple answer choices were deemed correct. To maintain the integrity of our evaluation, we removed questions deemed unanswerable or exhibiting multiple correct answers from the dataset. Questions with identifiable issues that could be rectified were corrected accordingly. This manual verification and refinement process removed a small percentage of questions originally collected for each dataset. Specifically, we removed 2.9% of questions from the MAET dataset, 3.5% from the CWET dataset, and 6.9% from the KCV dataset.
3.4. Benchmark Dataset and Evaluation
Based on the modeling and task descriptions of Section 3.1 and 3.2, we have created the following benchmark datasets:
Knowledge extraction dataset: As detailed in Section 3.2.1, these tasks are framed as MCQs derived from MITRE ATT&CK and CWE websites. We have released two datasets: MAET and CWET, comprising a total of 2036 MCQs. We show a sample in Table 1. Different LLMs are evaluated based on their accuracy in predicting the correct answers from the provided options, allowing us to measure their effectiveness in knowledge extraction.
Table 3. Sample from KCV and VOOD dataset
@@ 3.4. Benchmark Dataset and Evaluation
Knowledge understanding dataset: As explained in Section 3.2.2, we utilize CVE published in 2024 to create this boolean dataset. There are two variants: KCV, which includes the context from CVE JSON files, and VOOD, which lacks this context to assess the out-of-distribution performance of LLMs. Each dataset contains 466 Boolean questions. Along with the questions and answers, we also provide the JSON files used as context for evaluating the LLM’s comprehension abilities. A sample is shown in Table 3. Different LLM models are evaluated based on their accuracy in predicting the truthfulness of statements, both with and without context.
Table 4. Sample from CPST dataset
@@
Knowledge reasoning dataset: We have released RERT, a dataset consisting of 1000 questions based on security advisories from CISA, as detailed in Section 3.2.3. We show a sample in Table 2. Different LLM models are evaluated using the ROGUE-L metric (Lin, 2004) between the LLM-generated response and the ground truth. Additionally, we have released CPST, which includes 100 manually crafted CVSS3.1 vector strings along with their associated vulnerability scores. A sample of these is provided in Table 4. LLLM models are evaluated using the mean average deviation (MAD) against the ground-truth scores, assessing their accuracy in generating precise vulnerability assessments.
4. Experiments & Results
4.1. Generative LLMs
We evaluate 7 state-of-the-art LLMs varying in parameter size, organization, and access (See 5 for a comparative summary). We pick both open-source and closed models. Open-source models allow the download of full model weights whereas closed models provide an API for restricted access. We evaluate our benchmark against the following models:
Open-source models: Llama3-70B (lla, 2024a), Llama3-8B (lla, 2024b), Mistral-7B (Jiang et al., 2023), Mixtral-8x7b (mix, 2024)
Closed-source models: ChatGPT-3.5 (gpt, 2024a), ChatGPT-4 (gpt, 2024b), Gemini-Pro (v1.5) (gem, 2024)
Table 5. Comparision between different LLMs
Type Model Provider Training data # of parameters Context-window
Closed ChatGPT-3.5 OpenaAI Upto Sept. 2021 - 16385
ChatGPT-4 OpenaAI Upto Dec. 2023 - 128000
Gemini-Pro Google Upto Nov. 2023 - 128000
Open Llama3-70b Meta Upto Dec. 2023 70.6B 8000
Llama3-8b Meta Upto Mar. 2023 8B 8000
Mistral-7B Mistral Upto Sept. 2021 7B 8192
Mixtral-8x7B Mistral Upto Sept. 2021 46.7B 32000
@@ 4.1. Generative LLMs
Table 6. Result of different models on our benchmark tasks. $\uparrow$ and $\downarrow$ indicate higher and lower values are better.
4.2. Prompting Strategy for Evaluation
@@
We embedded relevant information from the JSON files directly within the prompt for the KCV task, which requires external knowledge. To maintain compatibility with model input constraints and avoid exceeding maximum token limits, we selectively included questions whose prompt length, including the embedded knowledge, did not exceed 5000 words. Although the models generally followed the specified output format, there were instances where responses deviated slightly from the instructions. In such cases, we manually corrected the outputs for accurate evaluation. We include all prompts we used to generate the responses in our dataset for reproducibility, and demonstrate more examples in Appendix B.
@@
4.3. Evaluation Metrics
We employ a range of evaluation metrics tailored to the specific nature of each task within our benchmarking framework:
  1. (1)
    MAET, CWET & KCV: For these three tasks, we utilize accuracy as the primary evaluation metric. Accuracy represents the percentage of questions answered correctly by the LLM.
  2. (2)
    VOOD: This out-of-distribution task presents models with True, False, and ”Don’t Know (X)” options. We consider a model’s prediction accurate if it selects ”Don’t Know (X),” demonstrating its ability to recognize when it lacks sufficient information to answer.
  3. (3)
    RERT: We treat the RERT task as a summarization task, evaluating generated summaries against reference summaries using the ROGUE-L metric (Lin, 2004). ROGUE-L measures the longest common subsequence (LCS) between the generated and reference summaries, capturing sentence-level similarity and coherence.
  4. (4)
    CPST: The CPST task requires models to predict a real number within the range of 0 to 10. We employ mean absolute deviation (MAD) to quantify prediction accuracy which calculates the average absolute difference between predicted and actual values.
4.4. Results Summary
Table 6 presents a comparative summary of the performance of various language models (LLMs) evaluated across the six benchmark tasks of SECURE. Our analysis reveals a consistent trend of closed-source models, particularly ChatGPT-4 and Gemini-Pro, exhibiting superior performance across most tasks.
@@ 4.4. Results Summary
5. Discussion & Analysis
5.1. Error Analysis
@@
The correct answer for this question is A) Sending legitimate control messages to devices. However, among all the LLMs evaluated, only ChatGPT-4 provided the correct response. This discrepancy might be due to the LLMs difficulty in recognizing the malicious intent embedded in the term ”legitimate control messages”. The positive sentiment associated with the word “legitimate” could have misled the models, causing them to overlook the actual malicious purpose.
@@
5.2. Impact of Confidence on LLM Accuracy
Prior studies have demonstrated that large language models (LLMs) can be calibrated through self-reflection or confidence analysis (Chen and Mueller, 2023). In this study, we evaluate the relationship between model confidence and performance across different confidence levels using the CWET task. Specifically, we measure the confidence of two representative LLMs, ChatGPT-4 and LLAMA3-70B, across five temperature settings: 0.6, 0.7, 0.8, 0.9, and 1.0, to obtain a more robust confidence estimate.
@@ 5.2. Impact of Confidence on LLM Accuracy
Figure 3. Confidence vs. Accuracy of LLMs
5.3. Open vs. Closed Model Performance
Figure 4. Performance Comparison of Open-Source and Closed-Source LLMs Across Different Tasks (Note: For CPST, lower scores indicate better performance).
@@
5.4. Evaluating Reasoning Abilities
Figure 5. Performance of LLMs when asked to perform step-by-step analysis
Figure 6. Inference time on reasoning tasks when asked to perform step-by-step analysis
@@ 5.4. Evaluating Reasoning Abilities
5.5. Variance in Prediction
Figure 7. Distribution of variance among predicitons for different LLMs
@@ 5.5. Variance in Prediction
5.6. Model Agreement Bias
@@
Figure 8. Fraction of incorrect prediction with no context and context
@@
Figure 8 illustrates the distribution of incorrect predictions for both tasks, categorized by whether the LLM agreed (predicted True) or disagreed (predicted False) with the statement. In VOOD, most errors stem from the LLMs agreeing with the statement, even though the statements related to vulnerabilities were not in their training data. This tendency to affirm novel information suggests a potential for hallucination, where the model generates plausible-sounding but unsubstantiated claims (Adlakha et al., 2024). It is also possible that the models are incorrectly associating information from known CVEs to newer, unseen vulnerabilities.
@@
Figure 9. Spearman correlation between different tasks
5.7. Task Correlation Analysis
@@
5.8. LLM Performance Across Expertise Levels
@@
Figure 10. Performance Comparison of LLMs across two different levels of question difficulty.
Figure 10 presents the average accuracy of all evaluated LLMs on the CWET and KCV tasks. Notably, LLMs demonstrate marginally superior performance on the “expert” questions for the CWET task, showing an improvement of 2.31% compared to the “novice” security questions. This suggests that LLMs can leverage their extensive training data to excel even in specialized cybersecurity domains for tasks primarily reliant on existing knowledge and pattern recognition. However, as discussed in Section 5.1, they often fail to retrieve the correct source for more factual questions. In contrast, the KCV task, which requires integrating new information and context-aware reasoning, reveals a different trend. Here, LLMs experience a significant decrease in accuracy, with a 5.44% drop for “expert” tasks compared to “novice” tasks. This finding underscores a potential limitation of LLMs: while proficient at utilizing pre-existing knowledge, their capability to effectively incorporate and reason with novel information, particularly in complex and dynamic fields like cybersecurity, necessitates further investigation and refinement.
5.9. Recommendations

### v3

@@ III-B1 Extraction Task
Use-case: Knowledge extraction tasks measure the ability of a language model to access its vast knowledge base and accurately recall specific facts. This is useful when a cybersecurity professional has to find answers to a threat, system weakness, mitigation plans, attack patterns, and vulnerability given a specific attack scenario or just simply to gain information to assist the other tasks they are performing. LLMs are considered to be repositories of world knowledge with few-shot learning [51] and zero-shot reasoning [52] skills and hence are frequently inquired to obtain answers for different questions. MAET and CWET, both extraction tasks, consist of such questions that replicate this situation and evaluate if LLMs have sufficient knowledge in the ICS domain. Since state-of-the-art LLMs are trained on world knowledge and have seen documents on MITRE and CWE, such questions measure whether they can accurately recall answers to these kinds of questions.
III-B2 Understanding Task
Our second task dataset is designed to evaluate the ability of LLM to comprehend and understand the security-related text. Given the continuously evolving landscape of cybersecurity, it is crucial that LLMs can assimilate new information and use it to generate accurate responses. For this purpose, we utilize the CVEs (Common Vulnerabilities and Exposures) published in 2024, available at CVE project repository [13]. CVE is better for this task since they are always updated but CWEs are more appropriate for the extraction task as they are updated less frequently.
To ensure the integrity of our evaluation, we verified that none of the pretrained models we selected for evaluation (ChatGPT-4 [1], ChatGPT-3.5 [18], Llama3-70b [19], Llama3-8b [20], Gemini-Pro [21], Mistral-7B [22] , Mixtral-8x7b [23]) had access to these CVEs during their training phase by confirming their training-cut-off date (See Table VIII). This precaution guarantees that the models have not been exposed to this specific information and thus must rely on their comprehension abilities, and therefore, maintain the validity and reliability of SECURE.
@@ III-B2 Understanding Task
Use-case: Knowledge understanding tasks are designed to assess the model’s capability to discern the truthfulness of statements with and without a context. KCV evaluates whether existing LLMs can comprehend security documents, and answer questions based on a given context. VOOD inspects how the models perform when the context is not provided. Both of these tasks measure the reliability of LLMs. An effective understanding of cybersecurity can only ensure that the LLM can accurately interpret and respond to nuanced security challenges.
TABLE II: An example row from RERT dataset. Overview & vulnerability description shortened for readability.
@@ 3.2.1 Path Traversal CWE-22 There are multiple ways in LAquis
III-B3 Reasoning Task
Our third task evaluates the LLM’s reasoning capability in the context of ICS security through the Risk Evaluation Reasoning Task (RERT). To create RERT, we compile a risk assessment dataset using cybersecurity advisories from the Cybersecurity and Infrastructure Security Agency (CISA) [15]. We focus on all significant ICS-related advisory reports, each containing an Executive Summary, Risk Evaluation, Technical Details (including Vulnerability Overview, Affected Products, Background Researcher), and Mitigations. Reports lacking essential information are discarded to ensure data quality, and the remaining comprehensive reports are processed for inclusion.
@@ III-B3 Reasoning Task
In addition, we create another dataset CPST (CVSS Problem Solving Task) to measure the problem-solving skills of LLMs in security. Specifically, we use the CVSS (Common Vulnerability Scoring System) framework and manually collect 100 unique CVSS3.1 vector strings from the Cybersecurity and Infrastructure Security Agency (CISA) [15]. These scores, in the range of 0-10, can be computed using the CVSS calculator which uses the Base, Temporal, and Environmental scores to determine the overall severity of a vulnerability [53]. This task involves using an existing formula from CVSS3.1 standard and computing a value. This computation allows the evaluation of the problem-solving skills of LLMs in practical security settings.
Use-case: LLMs are now integrated into emails and document software to summarize conversations and texts. Such LLMs work quite well in handling the general English language but summarizing a threat report is significantly different as the LLMs need to understand the technical details that consist of vulnerability, affected products, and risks. RERT dataset is designed to evaluate the ability of LLMs in summarizing an extensive threat report. The CVSS Problem Solving Task checks if LLMs can understand and use the CVSS formula without being explicitly programmed (zero-shot evaluation). While a simple program can do this deterministically, this task shows if LLMs can help in real-world scenarios where users might not know the formula. It shows LLMs’ ability to reason and solve problems, not just follow set rules.
III-C Dataset Validation
@@
Our analysis revealed several categories of issues within the generated questions. Some questions were deemed unanswerable from the provided context, particularly within the True/False format. In the Multiple Choice Question (MCQ) format, we encountered instances where multiple answer choices were deemed correct. To maintain the integrity of our evaluation, we removed questions deemed unanswerable or exhibiting multiple correct answers from the dataset. Questions with identifiable issues that could be rectified were corrected accordingly. We fixed 13 questions in MAET, 0 in CWE and 22 questions in KCV. This manual verification and refinement process removed a small percentage of questions originally collected for each dataset. Specifically, we removed 2.9% of questions from the MAET dataset, 3.5% from the CWET dataset, and 6.9% from the KCV dataset, totaling 29, 36, and 32 questions respectively.
III-D Benchmark Dataset and Evaluation
Based on the modeling and task descriptions of Section III-A and III-B, we have created the following benchmark datasets:
Knowledge extraction dataset: As detailed in Section III-B1, these tasks are framed as MCQs derived from MITRE ATT&CK and CWE websites. We have released two datasets: MAET and CWET, comprising a total of 2036 MCQs. We show a sample in Table I. Different LLMs are evaluated based on their accuracy in predicting the correct answers from the provided options, allowing us to measure their effectiveness in knowledge extraction.
TABLE III: Sample from KCV and VOOD dataset
@@ III-D Benchmark Dataset and Evaluation
Knowledge understanding dataset: As explained in Section III-B2, we utilize CVE published in 2024 to create this boolean dataset. There are two variants: KCV, which includes the context from CVE JSON files, and VOOD, which lacks this context to assess the out-of-distribution performance of LLMs. Each dataset contains 466 Boolean questions. Along with the questions and answers, we also provide the JSON files used as context for evaluating the LLM’s comprehension abilities. A sample is shown in Table III. Different LLM models are evaluated based on their accuracy in predicting the truthfulness of statements, both with and without context.
TABLE IV: Sample from CPST dataset
@@
Knowledge reasoning dataset: We have released RERT, a dataset consisting of 1000 questions based on security advisories from CISA, as detailed in Section III-B3. We show a sample in Table II. Different LLM models are evaluated using the ROGUE-L metric [54] between the LLM-generated response and the ground truth. Additionally, we have released CPST, which includes 100 manually crafted CVSS3.1 vector strings along with their associated vulnerability scores. A sample of these is provided in Table IV. LLM models are evaluated using the mean average deviation (MAD) against the ground-truth scores, assessing their accuracy in generating precise vulnerability assessments.
IV Experiments & Results
We evaluate 7 state-of-the-art LLMs varying in parameter size, organization, and access (See Appendix -A and Table VIII for detail and comparative summary). We pick both open-source and closed models based on LLM leaderboards (ChatArena [55]). Open-source models allow the download of full model weights whereas closed models provide an API for restricted access. We evaluate our benchmark against the following models:
Open-source models: Llama3-70B [19], Llama3-8B [20], Mistral-7B [22], Mixtral-8x7b [23]
Closed-source models: ChatGPT-3.5 [18], ChatGPT-4 [1], Gemini-Pro (v1.5) [21]
TABLE V: Result of different models on our benchmark tasks. $\uparrow$-$\downarrow$ indicate higher-lower values are better.
@@ IV Experiments & Results
IV-A Prompting Strategy for Evaluation
@@
We embedded relevant information from the JSON files directly within the prompt for the KCV task, which requires external knowledge. To maintain compatibility with model input constraints and avoid exceeding maximum token limits, we selectively included questions whose prompt length, including the embedded knowledge, did not exceed 5000 words. Although the models generally followed the specified output format, there were instances where responses deviated slightly from the instructions. In such cases, we manually corrected the outputs for accurate evaluation. We include all prompts we used to generate the responses in our dataset for reproducibility, and demonstrate more examples in Appendix -C.
@@
IV-B Evaluation Metrics
We employ a range of evaluation metrics tailored to the specific nature of each task within our benchmarking framework: accuracy for MAET, CWET, VOOD and KCV, ROGUE-L [54] for RERT and mean absolute deviation (MAD) for CPST. Because of the page-limit, we explain the metrics in detail in Appendix -B.
IV-C Results Summary
Table V presents a comparative summary of the performance of various language models (LLMs) evaluated across the six benchmark tasks of SECURE. Our analysis reveals a consistent trend of closed-source models, particularly ChatGPT-4 and Gemini-Pro, exhibiting superior performance across most tasks.
@@ IV-C Results Summary
V Discussion & Analysis
V-A Error Analysis
@@
The correct answer for this question is A) Sending legitimate control messages to devices. However, among all the LLMs evaluated, only ChatGPT-4 provided the correct response. This discrepancy might be due to the LLMs difficulty in recognizing the malicious intent embedded in the term “legitimate control messages”. The positive sentiment associated with the word “legitimate” could have misled the models, causing them to overlook the actual malicious purpose.
@@
V-B Impact of Confidence on LLM Accuracy
Prior studies have demonstrated that large language models (LLMs) can be calibrated through self-reflection or confidence analysis [56]. In this study, we evaluate the relationship between model confidence and performance across different confidence levels using the CWET task. Specifically, we measure the confidence of two representative LLMs, ChatGPT-4 and Llama3-70B, across five temperature settings: 0.6, 0.7, 0.8, 0.9, and 1.0, to obtain a more robust confidence estimate.
@@ V-B Impact of Confidence on LLM Accuracy
Figure 3: Confidence vs. Accuracy of LLMs
V-C Open vs. Closed Model Performance
@@
Figure 4: Performance Comparison of Open-Source and Closed-Source LLMs Across Different Tasks (Note: For CPST, lower scores indicate better performance).
@@
V-D Evaluating Reasoning Abilities
Figure 5: Performance of LLMs when asked to perform step-by-step analysis
Figure 6: Inference time on reasoning tasks when asked to perform step-by-step analysis
@@ V-D Evaluating Reasoning Abilities
V-E Variance in Prediction
Figure 7: Distribution of variance among predictions for different LLMs
@@ V-E Variance in Prediction
V-F Model Agreement Bias
@@
Figure 8: Fraction of incorrect prediction with no context and context
@@
Figure 8 illustrates the distribution of incorrect predictions for both tasks, categorized by whether the LLM agreed (predicted True) or disagreed (predicted False) with the statement. In VOOD, most errors stem from the LLMs agreeing with the statement, even though the statements related to vulnerabilities were not in their training data. This tendency to affirm novel information suggests a potential for hallucination, where the model generates plausible-sounding but unsubstantiated claims [57]. It is also possible that the models are incorrectly associating information from known CVEs to newer, unseen vulnerabilities.
@@
V-G Task Correlation Analysis
@@
Figure 9: Spearman correlation between different tasks in SECURE benchmark
@@
V-H LLM Performance Across Expertise Levels
Figure 10: Performance comparison of LLMs across two different levels of question difficulty.
@@ V-H LLM Performance Across Expertise Levels
Figure 10 presents the average accuracy of all evaluated LLMs on the CWET and KCV tasks. Notably, LLMs demonstrate marginally superior performance on the “expert” questions for the CWET task, showing an improvement of 2.31% compared to the “novice” security questions. This suggests that LLMs can leverage their extensive training data to excel even in specialized cybersecurity domains for tasks primarily reliant on existing knowledge and pattern recognition. However, as discussed in Section V-A, they often fail to retrieve the correct source for more factual questions.
In contrast, the KCV task, which requires integrating new information and context-aware reasoning, reveals a different trend. Here, LLMs experience a significant decrease in accuracy, with a 5.44% drop for “expert” tasks compared to “novice” tasks. This finding underscores a potential limitation of LLMs: while proficient at utilizing pre-existing knowledge, their capability to effectively incorporate and reason with novel information, particularly in complex and dynamic fields like cybersecurity, necessitates further investigation and refinement.
V-I Human Benchmark
We conducted an experiment involving human subjects where participants were asked to answer a subset of the SECURE benchmark questions. 7 PhD students and 7 industry professionals working in cybersecurity and threat intelligence with years of experience ranging from 2-10 years were participants of this study.
We provided an MCQ form to each participant consisting of 75 questions from MAET, CWET, and KCV datasets divided into three sets, and were asked to provide their answers. We also collected their confidence level (high or low) when providing the answers and the total time taken for each set. Table VI shows the average performance of the human subjects on the dataset. Comparing the results with Table V, we can observe that large language models outperform human subjects on the three tasks. Most participants considered KCV a difficult task as it involves CVE document analysis before validating or refuting the given statement. On average, PhD students scored higher accuracy than industry professionals. While analysts scored an average accuracy of 72.2% accuracy on all tasks, PhD students scored 77.3% accuracy.
TABLE VI: Average performance of human subjects on the subset of SECURE benchmark
Dataset Accuracy (Average) Confidence (Average) Average time (mins)
MAET 77% 76% 25
CWET 78% 81% 23
KCV 68% 72% 30
V-J Recommendations